# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ 4.1 Multitask Learning Benchmarks
4.1.1 Performance Evaluation on Atari 100k benchmark
Setup. We evaluate ScaleZero on the large-scale Atari 100k benchmark, which comprises 26 distinct games. This benchmark is renowned for its significant heterogeneity across tasks, featuring diverse observation spaces (pixel inputs), action spaces (discrete instruction sets), and widely varying game difficulties and mechanics. We compare the multitask (MT) ScaleZero model against a single-task (ST) UniZero baseline, which is individually trained and optimized for each game. The evaluation metrics are the mean and median of the Human-Normalized Score (HNS), a standard in the field, to provide a comprehensive assessment of performance across tasks of varying difficulty.
Results. As shown in Table 1, by training a single, unified model, ScaleZero achieves a mean normalized score that surpasses the average performance of 26 independently trained single-task UniZero models. This indicates that ScaleZero not only successfully learns a diverse set of tasks but also achieves positive knowledge transfer across them. Notably, the performance gains are most pronounced on several notoriously challenging tasks, such as Seaquest. We hypothesize that this is because ScaleZero’s unified world model can learn general visual and dynamic priors from simpler tasks and effectively apply them to more difficult ones that demand complex planning and exploration strategies. Although the median score is slightly lower than the single-task baseline, which is primarily influenced by a few tasks with extremely hard exploration challenges, the superior mean score robustly demonstrates the overall advantage of our approach.
Table 1: Comparison of Human-Normalized Scores on 26 Atari games for ScaleZero (MT) vs. UniZero (ST). Detailed curves are available in Appendix B Figure 9.
Algorithm Normalized Mean Normalized Median
@@ 4.1.1 Performance Evaluation on Atari 100k benchmark
4.1.2 Performance Evaluation on the DeepMind Control Suite
Setup. To test the generalization capability of ScaleZero in the continuous control domain, we conducted evaluations on the DMC benchmark, which includes 18 tasks based on the MuJoCo physics engine. These tasks feature vector state space and continuous action spaces, posing higher demands on the model’s fine-grained control and dynamics modeling abilities. ScaleZero also leverages the principles of Sampled Policy Iteration (Hubert et al., 2021) same as UniZero. Note that because observation space of DMC is vector, we use the MLP encoder. The experimental setup mirrors that of the Atari experiments, comparing the multitask ScaleZero against 18 independent single-task UniZero models.
Results. The results in Table 2 show that ScaleZero achieves returns comparable to or even higher than the single-task baselines on the vast majority of DMC tasks. The learning curves, as illustrated in Figure 10, clearly reveal that ScaleZero can effectively handle the diverse physical dynamics and control requirements of different tasks within a single, unified model. This success validates the versatility of the ScaleZero architecture, proving its competitiveness in complex control problems involving high-dimensional continuous states and actions.
@@ 4.1.2 Performance Evaluation on the DeepMind Control Suite
Table 2: Performance comparison between ScaleZero (MT) and UniZero (ST) across various tasks in DMControl. For reference, L2M (MT) Schmied et al. (2023) obtained a mean score of 840 under an offline supervised learning from expert data setting. The best performance for each task is indicated in bold. ScaleZero’s higher human-normalized median scores highlight its strong performance across continuous action spaces. Detailed curves are available in Appendix C  Figure 10.
4.1.3 Performance Evaluation on the Jericho Text-Based Adventure Games
Setup. To further evaluate ScaleZero’s text-based reasoning and planning, we assess it on the Jericho (Hausknecht et al., 2020) benchmark of interactive fiction, where both observations and actions are expressed in natural language. Jericho features a combinatorial action space and sparse, delayed rewards, and it demands strong long-horizon contextual reasoning. We use four representative games—Acorncourt, Omniquest, Detective, and Zork1—covering diverse narrative structures and interaction complexity. For parity, both the single-task UniZero and the multitask ScaleZero employ the same BGE text encoder (Xiao et al., 2023). In addition, we compare against CALM+OC, a recent LM-in-the-Loop method that adaptively updates the action generator using state-feature–based categorization of gameplay transitions (Sudhakar et al., 2023). The primary evaluation metric is the average return measured after training convergence.
Results. The results in Table 3 show that ScaleZero achieves returns comparable to the single-task UniZero baseline and generally outperforms the CAML+OC method across most games. The learning curves in Figure 13 further indicate that a single, unified model can efficiently handle Jericho’s combinatorial language action space and sparse, delayed rewards. These findings validate the applicability of the ScaleZero architecture for text-based RL, demonstrating competitive performance on tasks that require language understanding and long-horizon planning.
Table 3: Average returns of ScaleZero and other baselines on four Jericho tasks. CALM+OC(Sudhakar et al., 2023) denotes the variant of CALM with online categorization based on state features. Detailed learning curves for ScaleZero and UniZero are provided in Appendix D  Figure 13.
@@ 4.2 Efficiency Evaluation of Dynamic Parameter Scaling
This section aims to evaluate the effectiveness of our LoRA-based online dynamic parameter scaling strategy integrated into ScaleZero, termed ScaleZero-DPS, in improving sample and computational efficiency. To this end, we conducted a comparative analysis between the standard ScaleZero and ScaleZero-DPS on the full DMC18 benchmark. The core evaluation metric is defined as the number of environment interaction steps required to reach a performance level comparable to the single-task baseline (i.e., 100% of its performance). As illustrated in Figure 5, ScaleZero-DPS achieves the target performance using only about 80% of the interaction steps required by the standard ScaleZero. This result highlights the significant efficiency gains afforded by our dynamic scaling strategy. Furthermore, the detailed performance curves, presented in Appendix Figure 11, reveal that ScaleZero-DPS exhibits a steeper learning curve. This indicates a faster learning rate, leading to higher sample and computational efficiency throughout the training process.
Figure 5: Comparison of interaction cost for ScaleZero vs. ScaleZero-DPS on DMC18. ScaleZero-DPS reaches the target performance with approximately 80% of the interaction steps. For detailed performance curves, please refer to Appendix C Figure 11.
@@ 4.3 Understanding MoE Efficacy: An Empirical and Theoretical Analysis
The demonstrated success of Mixture-of-Experts (MoE) in multi-task scenarios motivates a deeper investigation into its underlying mechanisms. In the following, we explore these mechanisms from both experimental and theoretical perspectives. We begin by summarizing our empirical findings, which indicate that replacing standard MLP layers with MoE counterparts significantly reduces gradient conflicts. Notably, this reduction is not confined to the MoE layers themselves but also extends to adjacent network components, such as their input (detailed experimental settings and results are provided in Section˜E.1).
To complement our empirical analysis, we provide a theoretical justification for how MoE mitigates gradient conflict. Our theory builds upon prior work (Chen et al., ) that characterizes MoE training by two distinct phases: an initial exploration stage, where inputs are routed nearly uniformly across experts, followed by a router learning stage, where the router gradually specializes in directing inputs. Grounded in this two-stage framework, we propose an informal theorem asserting that, regardless of the training stage, the upper bound on gradient conflict for an MoE layer is strictly lower than that of its MLP counterpart, and that this reduction is primarily governed by the router’s behavior.
@@
In a Mixture-of-Experts (MoE) layer, consider two tasks $t_{1}$ and $t_{2}$ with routing weights $\lambda^{t_{1}}_{m},\lambda^{t_{2}}_{m}$ over $M$ experts, and per-task gradients on expert $m$, denoted as $g_{t_{1}}^{(m)}$ and $g_{t_{2}}^{(m)}$. Let $G$ be the maximum gradient conflict on any single expert, and $\cos(u,v)$ denote the cosine similarity between vectors $u$ and $v$. Then the following results hold:
(1). (Generally) Define the $m$-th element of vectors $u,v$ as $u_{m}=\lambda^{t_{1}}_{m}\|g_{t_{1}}^{(m)}\|,\;v_{m}=\lambda^{t_{2}}_{m}\|g_{t_{2}}^{(m)}\|,\;m=1,\dots,M.$ The full-layer gradient conflict admits a general upper bound: $\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\cos(u,v)$.
(2). (Sparse Router) In exploration stage, where tasks uniformly select experts, the expected gradient conflict of the MoE layer is approximately bounded by $G\cdot(1-M!/((M-K)!\,M^{K})$, where $K$ is the task number.
(3). (Sparse Router) In the router learning stage, where the expert sets selected by the router gradually stabilize, the gradient conflict of the MoE layer is roughly upper bounded by $\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}$, where $a=|S_{i}|$ and $b=|S_{j}|$ denote the sizes of the expert sets selected by tasks $i$ and $j$, and $U=|S_{i}\cap S_{j}|$ denotes the size of their intersection.
In Section˜E.2, we provide detailed explanations of the the three parts of this theorem and their related corollaries, along with complete proofs.
@@ 5 Related Work
MCTS-based Planning with Learned World Models. The integration of MCTS with deep neural networks has become a cornerstone of modern reinforcement learning, particularly in domains requiring sophisticated planning. The seminal AlphaZero series (Silver et al., 2016; 2017) established a powerful paradigm by coupling MCTS with learned policy and value networks. This approach was further advanced by MuZero (Schrittwieser et al., 2019), which obviated the need for an explicit environment simulator by learning a latent world model to jointly predict future rewards, policies, and values. This architecture enables planning entirely within a learned representation space. Subsequent research has extended the MuZero framework, focusing on enhancing model capabilities (Hubert et al., 2021; Antonoglou et al., 2021), improving sample efficiency (Danihelka et al., 2022; Ye et al., 2021), and refining training stability (Schrittwieser et al., 2021). Notably, the advent of Transformer-based world models (Micheli et al., 2022; Robine et al., 2023), as exemplified by UniZero (Pu et al., 2024), has significantly augmented representational capacity, enabling stronger long-horizon planning and generalization. However, the monolithic, dense representations that enable single-task excellence become a liability in multi-task settings, causing representational interference and destructive interference. Simply scaling these models exacerbates this architectural bottleneck. Our work try to confront this fundamental limitation.
Multi-Task Reinforcement Learning. To address the challenge of generalization across diverse tasks, MTRL aims to leverage shared knowledge to improve data efficiency and performance(Vithayathil Varghese & Mahmoud, 2020; D’Eramo et al., 2024). A prevalent architectural motif is the shared-backbone with task-specific heads (Kumar et al., 2022; Hansen et al., 2023; 2022; Lee et al., 2022a; Gallouédec et al., 2024a). While promoting knowledge transfer, this design is insufficient for world models where the core challenge lies in disentangling the latent dynamics predictions within the backbone itself, not just the final policy or value outputs. Recent research in MTRL has pursued several avenues to mitigate these issues. From a representational standpoint, context-based methods(Rakelly et al., 2019; Sodhani et al., 2021; Seo et al., 2022) condition the model on task metadata to enable selective knowledge sharing, while others introduce learnable modulation modules to preserve task-specific knowledge (Schmied et al., 2023). Architecturally, modular designs have emerged as a promising direction(Lan et al., 2023; Yang et al., 2020; Sun et al., 2022). These include soft modularization with task-aware routing (Yang et al., 2020; He et al., 2023; Hendawy et al., 2023) and constructing policy subspaces via parameter composition (Sun et al., 2022; 2023). On the optimization front, various strategies(Fernando et al., 2023; Lin et al., 2024; Ma et al., 2023) have been developed to manage conflicting gradients, a critical impediment in MTRL. These include projecting conflicting gradients into a non-conflicting space (Fernando et al., 2023) and harmonizing training through dynamic loss and gradient normalization (Lin et al., 2024; Ma et al., 2023; Chen et al., 2018). While effective at mitigating conflicts, these methods primarily address the symptoms (conflicting gradients) rather than the root cause: the architectural bottleneck of a model lacking sufficient specialized capacity to represent diverse task dynamics. These diverse approaches underscore the central challenge in MTRL: balancing knowledge sharing with task specialization. Our work seeks to achieve this balance through a novel architectural design.
Sparse and Parameter-Efficient Architectures. Scaling deep RL models is computationally expensive (Espeholt et al., 2018). Sparse activation models, like the Mixture-of-Experts (MoE) (Dai et al., 2024), address this by increasing representational power at a near-constant computational cost via conditional computation. Their success in single-task RL (Obando-Ceron et al., 2024) highlights their potential as a native architectural prior for multi-task specialization. Concurrently, parameter-efficient fine-tuning (PEFT) methods (Dou et al., 2024; Dettmers et al., 2023; Hayou et al., 2024; Liu et al., 2024c) like Low-Rank Adaptation (LoRA) (Hu et al., 2022) enable lightweight, task-specific adaptations by tuning a small subset of parameters, avoiding the cost of full fine-tuning in MTRL (Wang et al., 2023; Yang et al., 2025; Agiza et al., 2024; Zhang et al., 2025). We identify a critical synergy between these two approaches. MoE enables coarse-grained, input-dependent specialization by routing computations to expert sub-networks, tackling representational entanglement at an architectural level. In parallel, LoRA facilitates fine-grained, task-specific adaptation in a parameter-efficient manner, injecting nuanced adjustments without catastrophic interference. The synergistic integration of MoE’s scalable representation with LoRA’s efficient adaptation is a critical frontier, yet one that has seen little exploration(Dou et al., 2023). Our work bridges this gap by proposing a novel architecture that unifies MoE and LoRA within a Transformer-based world model.
@@ 6 Conclusion and Future work
In this paper, we examine the diminishing plasticity phenomenon in unified world models under multi-task scenarios and quantify its impact on model performance. By exploring the expansive design space defined by the input–model–optimization axes, we introduce ScaleZero—a principled method that carefully balances performance with sample efficiency. Our extensive experiments on pure online RL benchmarks—which include 26 Atari, 18 DMControl, and 4 Jericho environments—demonstrate that ScaleZero achieves average returns comparable to or even surpassing those of specialized single-task baselines. To further balance computational costs, we introduced a LoRA-based dynamic parameter scaling (DPS) strategy. This strategy has currently only been prototyped on the DMC benchmark, where it achieved competitive performance with 80% of the interaction steps, showcasing its potential for improving sample efficiency. Despite these promising results, several avenues remain open for future investigation: A comprehensive evaluation of the DPS strategy’s generalization, scalability, and overhead is required on broader task sets like Atari. Investigating deeper synergistic mechanisms between MoE and LoRA to further enhance model expressivity and adaptability. Combining our online framework with offline pre-training of large-scale models to potentially boost both efficiency and final performance. We believe the continued exploration of these directions will advance multi-task reinforcement learning toward the goal of general intelligence.

### v2

@@ 4.1 Multitask Learning Benchmarks
4.1.1 Atari Multitask Benchmark
Table 1: Performance comparison of our multi-task model, ScaleZero (MT), against the single-task UniZero (ST) baseline across discrete (Atari) and continuous (DMControl) control benchmarks.
Algorithm Norm. Mean Norm. Median
@@ 4.1.1 Atari Multitask Benchmark
(a) Human-normalized scores on 26 Atari games. ScaleZero achieves a higher mean score. Full curves are in Appendix B. We omit other MT baselines as we are unaware of prior work tackling all 26 games with a single online RL agent.
@@
(b) Raw scores on 18 DMControl tasks. ScaleZero achieves a superior median score, indicating robust generalist performance. Full curves are in Appendix C. For reference, L2M (MT) Schmied et al. (2023) reported a mean score of 840 on a 10-task subset of this benchmark, trained via offline supervised learning from expert data.
Setup. We evaluate ScaleZero on the 26 games of the Atari 100k benchmark, a standard testbed for heterogeneous MTRL due to its diverse visual styles, action spaces, and game mechanics. We compare the single multi-task (MT) ScaleZero agent against a strong single-task (ST) UniZero baseline, where 26 separate models are trained for each game. Performance is measured using the standard Human-Normalized Score (HNS) (Agarwal et al., 2021), reporting both mean and median to capture overall capability. Experimental details are provided in Appendix B.
Results. As shown in Table 1, ScaleZero, as a single generalist agent, achieves a mean normalized score that surpasses the performance of 26 independently trained single-task specialists. This result demonstrates that ScaleZero effectively mitigates the negative interference observed in heterogeneous MTRL, enabling robust knowledge transfer. Notably, performance gains are particularly pronounced on notoriously difficult exploration tasks like Seaquest—a game we identified as a primary failure case for baseline models in our initial diagnosis (Figure 1) While the median score is slightly lower, influenced by a few hard-exploration games, the higher mean score confirms the overall advantage and generalization of our approach.
4.1.2 DeepMind Control Suite
Setup. To test ScaleZero’s generalization to continuous control, we evaluate it on 18 tasks from the DeepMind Control (DMC) Suite. These tasks, based on the MuJoCo physics engine (Todorov et al., 2012), feature low-dimensional state inputs and continuous action spaces, demanding precise dynamics modeling. Since observations are vector-based, we use an MLP encoder. The experimental protocol remains consistent, comparing the single MT ScaleZero agent against 18 individually trained ST UniZero models. Experimental details are provided in Appendix C.
Results. 1(b) demonstrate that ScaleZero achieves competitive performance, surpassing the baseline on several tasks and achieving a superior median score. This metric is particularly indicative of a strong generalist agent, as it shows robust performance across the majority of tasks rather than excelling on a few easy ones. This success validates the versatility of the ScaleZero, proving its effectiveness extends beyond visual domains to distinct and complex continuous control problems.
4.1.3 Jericho Text-Based Adventure Games
Setup. To assess performance on tasks requiring language understanding and long-horizon planning, we use the Jericho benchmark. These games feature combinatorial action spaces and sparse rewards, posing a reasoning challenge. We evaluate on four representative games and compare against the ST UniZero baseline and CALM+OC (Sudhakar et al., 2023), a recent language-model-based method. Both ScaleZero and UniZero use the same BGE text encoder (Xiao et al., 2023) for fair comparison.
Results. As reported in Table 2, ScaleZero achieves performance on par with specialized single-task agents and is competitive with the strong CALM+OC baseline. This shows that a unified ScaleZero agent can effectively learn policies in environments where both states and actions are linguistic. These findings confirm that the design of ScaleZero are modality-agnostic, successfully extending to the domain of text-based reasoning and planning. Experimental details are provided in Appendix D.
Table 2: Average returns comparison on four Jericho tasks. Detailed learning curves are in Appendix D.
@@ 4.2 Efficiency Evaluation of Dynamic Parameter Scaling
Having established the performance of ScaleZero, we now evaluate the efficiency gains from our Dynamic Parameter Scaling strategy. We compare the standard ScaleZero against ScaleZero-DPS on the DMC18 benchmark. The primary metric is the number of environment interactions required to match the final performance of the single-task baselines. As shown in Figure 4, ScaleZero-DPS achieves this performance target using only 71.5% of the interactions needed by the standard ScaleZero. This corresponds to a substantial 28.5% reduction in sample complexity, directly validating the effectiveness in resolving the static resource allocation problem. The steeper learning curves shown in Figure 11 further confirm that this gain is consistent throughout the training process. A mechanistic analysis in Appendix C.4 reveals the agent’s efficiency stems from an emergent hierarchical strategy. The model learns to preserve foundational knowledge in early layers while applying targeted plasticity in later layers to adapt to new tasks, providing a clear mechanistic explanation for its performance.
Figure 4: Comparison of interaction cost for ScaleZero vs. ScaleZero-DPS on DMC18. ScaleZero-DPS reaches the target performance with a 28.5% reduction in environment steps. Detailed curves are in Appendix C.
@@ 4.3 Understanding MoE Efficacy: An Empirical and Theoretical Analysis
The superior performance of ScaleZero motivates a deeper investigation into its core architectural component: the Mixture-of-Experts backbone. Here, we analyze the mechanisms of MoE from both empirical and theoretical perspectives. Our empirical findings (detailed in Section˜E.1) reveal that replacing dense MLP layers with MoE counterparts not only reduces gradient conflict within the MoE layers themselves but also in adjacent network components. To explain this phenomenon, we provide a theoretical justification grounded in the two-phase training characterization of MoE (Chen et al., 2022). We propose the following informal theorem, which asserts that an MoE layer maintains a strictly lower upper bound on gradient conflict than a dense equivalent, governed primarily by the router’s specialization. Detailed explanations, proofs and corollaries are provided in Section˜E.2.
@@
In a Mixture-of-Experts (MoE) layer, consider two tasks $t_{1}$ and $t_{2}$ with routing weights $\lambda^{t_{1}}_{m},\lambda^{t_{2}}_{m}$ over $M$ experts, and per-task gradients on expert $m$, denoted as $g_{t_{1}}^{(m)}$ and $g_{t_{2}}^{(m)}$. Let $G$ be the maximum gradient conflict on any single expert, and $\cos(u,v)$ denote the cosine similarity between vectors $u$ and $v$. Then the following results hold: (1). (Generally) Define the $m$-th element of vectors $u,v$ as $u_{m}=\lambda^{t_{1}}_{m}\|g_{t_{1}}^{(m)}\|,\;v_{m}=\lambda^{t_{2}}_{m}\|g_{t_{2}}^{(m)}\|,\;m=1,\dots,M.$ The full-layer gradient conflict admits a general upper bound: $\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\cos(u,v)$. (2). (Sparse Router) In the exploration stage, where tasks uniformly select experts, the expected gradient conflict of the MoE layer is approximately bounded by $G\cdot(1-M!/((M-K)!\,M^{K}))$, where $K$ is the number of tasks. (3). (Sparse Router) In the router learning stage, where expert sets stabilize, the gradient conflict is roughly upper bounded by $\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}$, where $a=|S_{i}|$, $b=|S_{j}|$ are the sizes of the expert sets for tasks $i$ and $j$, and $U=|S_{i}\cap S_{j}|$ is their intersection size.
@@ 5 Related Work
MCTS with Learned World Models. Planning in a learned latent space, popularized by MuZero (Schrittwieser et al., 2019), is a dominant RL paradigm (Hubert et al., 2021; Antonoglou et al., 2021; Danihelka et al., 2022; Xuan et al., 2024; Niu et al., 2024). Recent works have integrated Transformers to enhance representational capacity (Micheli et al., 2022; Pu et al., 2024; Zhou et al., 2024), achieving state-of-the-art performance in single-task domains. However, their monolithic design is a critical liability in heterogeneous multitask settings, where they suffer from representational interference and plasticity collapse. Our work directly confronts this architectural bottleneck.
Multitask Reinforcement Learning. MTRL aims to improve data efficiency by sharing knowledge across tasks (Vithayathil Varghese & Mahmoud, 2020; Kumar et al., 2022; Hansen et al., 2023; Lee et al., 2022; Gallouédec et al., 2024). Existing methods mitigate interference through architectural solutions like task-specific modules (Schmied et al., 2023; Sun et al., 2022) or optimization-based approaches that manage gradient conflicts (Fernando et al., 2023; Ma et al., 2023). These strategies, however, have been predominantly studied outside of latent-space planning, where the unique challenge of disentangling dynamics prediction remains largely unaddressed.
Sparse and Parameter-Efficient Architectures. We draw inspiration from two complementary paradigms. Sparse MoE models offer a natural architectural prior for multitask specialization by enabling conditional computation (Dai et al., 2024; Obando-Ceron et al., 2024). Concurrently, parameter-efficient methods like LoRA (Hu et al., 2022) provide a lightweight mechanism for task-specific adaptation (Wang et al., 2023; Zhang et al., 2025). The synergy of integrating these paradigms within a transformer-based world model is a promising but underexplored frontier (Dou et al., 2023; Huang et al., 2024) that we bridge in this work. See Appendix F for more discussion.
@@ 6 Conclusion and Future work
In this work, we diagnose and addressed the critical issue of plasticity collapse in multitask world models. We introduced ScaleZero, a unified architecture incorporating a sparse MoE backbone to create specialized computational pathways. Our experiments across 48 distinct tasks (Atari, DMControl, Jericho) demonstrate that ScaleZero robustly matches or surpasses the performance of specialized single-task agents. Furthermore, we propose DPS, a LoRA-based strategy that achieves competitive performance on DMControl while using just 71.5% of the environment interactions, highlighting a promising path toward greater sample efficiency. Our work provides a robust architectural foundation for developing more capable and efficient generalist agents. See Appendix G for a full discussion.
Ethics Statement
We have read and adhered to the ICLR Code of Ethics. Our research focuses on the development of novel algorithms for multitask reinforcement learning, specifically proposing ScaleZero and the Dynamic Parameter Scaling (DPS) strategy. All experiments were conducted on standard, publicly available benchmarks (Atari, DMC, and Jericho), which are simulated environments. Our work does not involve human subjects, nor does it use any personally identifiable or sensitive data. The proposed methods are foundational and do not have immediate, direct real-world applications that could pose ethical risks or societal harm. To the best of our knowledge, our research does not create or exacerbate discrimination or bias, as the training environments do not contain demographic or social data. We have strived to uphold the principles of scientific integrity and transparency by providing a thorough description of our methods and experimental setup. We declare no conflicts of interest.
Reproducibility Statement
To ensure the reproducibility of our findings, we provide comprehensive details in the appendix of this paper. The appendix includes: (1) A detailed description of the implementation of our proposed model, ScaleZero, and the online Dynamic Parameter Scaling (DPS) strategy. (2) A thorough account of the experimental setup, including all hyperparameters, evaluation protocols, and the computational infrastructure used. (3) The complete theoretical proofs for the claims made in the paper. Furthermore, we commit to releasing our source code publicly upon the completion of the peer-review process to allow for full verification of our results and to facilitate future research in this area.
Appendix Contents
  • A.
    Implementation Details ........................................................................................................................................................................A
  • B.
    Atari Experiment Details ........................................................................................................................................................................B
  • C.
    DMC Experiment Details ........................................................................................................................................................................C
  • D.
    Jericho Experiment Details ........................................................................................................................................................................D
  • E.
    Gradient Analysis in MoE ........................................................................................................................................................................E
  • F.
    Detailed Related Work ........................................................................................................................................................................F
  • G.
    Detailed Conclusion and Future Work ........................................................................................................................................................................G
  • H.
    The Use of Large Language Models ........................................................................................................................................................................H